answersLogoWhite

0

🎒

Probability

The probability of a certain event is a number expressing the likelihood that a specific event will occur, expressed as the ratio of the number of actual occurrences to the number of possible occurrences. In mathematics, it is a measure of how often an event will happen and is the basis of statistics.

14,643 Questions

Can a probability be -0.0001?

No.

All probabilities lie between 0 and 1 inclusive; -0.0001 is not in that range.

How many different amounts can you have with only 2 coins?

That depends what coins are allowed. Since I don't know in what country you live, there is really no way to know. For "n" different coins, you have "n" options where the same coin is used twice; you also have n(n-1)/2 options that use two different coins.

What is the probability for rolling a number cube?

For most people the probability is 1: they have already rolled a number cube. For infants, it is quite likely to happen in the course of their lives, so again the probability is very close to 1.

How many permutations are in the word October?

There are 7 factorial, or 5,040 permutations of the letters of OCTOBER. However, only 2,520 of them are distinguishable because of the duplicate O's.

What is direct sampling?

When you taste the wine first secretly before buying at the market.

Are you more likely to get hit by lightning than win the lottery?

The odds of being struck by lightning are 576,000 to 1.

The odds of being killed by lightning are2,320,000 to 1.

The odds of winning the UK lotto are 1 in 13,983,816 (approx 1 in 14million).

So, you're more likely to get hit by lightning than win the lottery.

When is a one sample t-test used?

When the hypotheses being tested is one sided. That is to say, the hypothesis is not that the random variable X is different to some value V, but either that X > V or X < V. Of course, either hypothesis could include equality.

When 2 coins are tossed simultaneously what is probabiliy of getting atleast two heads?

At least two heads with two coins? You can't get more.

There are 4 different outcomes:

tail-tail, head-tail, tail-head and head-head.

You can use one out of four - which gives us the probability 1/4 = 0.25 = 25%

A whole number consist ofthe digits 24680 how many whole numbers are possible?

There are 5 digits in 24680; think of them as 5 compartments. You can select 5 numbers for the 1st compartment, 4 numbers for the 2nd compartment, 3 numbers for the 3rd compartment, 2 numbers for the 4th compartment, and 1 number for the 5th compartment. So, the answer is 5*4*3*2*1 or 5! or 120.

What does DC mean in the event of a miscarriage?

I think you may be referring to "D and C," which means "Dilation and Curettage." It is a procedure to scrape and collect the tissue (endometrium) from inside the uterus. Dilation is a widening of the cervix to allow instruments into the uterus and curettage is the scraping of the contents of the uterus. The procedure has a number of uses including the removal of tissue after a miscarriage. For more information, go to: http://www.emedicinehealth.com/articles/12190-1.asp

If two fair dice were rolled what is the probability of them rolling a sum of three?

Each cube can come up in 6 different ways.

Total possible outcomes for 2 cubes is (6 x 6) = 36 .

There are 2 ways to roll a sum of three:

1, 2

2, 1

The probability of success is (2/36) = 5 and 5/9 percent

What are the odds of rolling either a 1 or a 5 with 3 dice?

well there are 6 sides and the odd of hitting one

is a 1 out of 6 chance. the odds of hitting five is a 5 out 6

chance

An expression of the risk associated with a hazard that combines its severity and probability into a single Arabic numeral which can be used to help determine hazard abatement priorities.?

The combination (product, actually) of a probability and the associated benefit (or cost) of a certain situation is called the "mathematical expectation", "expectation value", or "expected value".

What is the chance of rolling an even number on both of the 2 dice?

There are four outcomes possible. Both even, both odd, and one of each twice. So, in one roll, it looks like your chances are 1 in 4. Is that what you are after?